OpenStudy (fairegaming):

I need help with Graphing Quadratic Functions - *MEDAL*! ---------------------------------------------------------------------- Just show me how and answer a few. --------------------------------------------- 1.) y = x(2nd) -6x + 4 2.) y = 1/2x(2nd) + 6 3.) y = -x(2nd) + 4x - 4 4.) y + 4 = 3x(2nd) -Just incase the (2nd) means (Power to the 2nd)-

OpenStudy (fairegaming):

hello?

alones (alones):

So you need each of them answered?

OpenStudy (fairegaming):

I just need the steps but if you can do one and show me how it would be awesome.

OpenStudy (fairegaming):

not trying to make you do all the work :P

alones (alones):

Mkay u mean y = x\(^{2}\)-6x + 4

OpenStudy (fairegaming):

yes

alones (alones):

Okay that makes more sence, so u need this graphed?

OpenStudy (fairegaming):

Yup

OpenStudy (rebecca):

https://www.desmos.com/calculator

alones (alones):

Follow beccas steps lol plug them in and it will hsot u where it's graphed ;)

OpenStudy (fairegaming):

its not just that. I have to find the vertex and etc. until I find the point

OpenStudy (calculusxy):

Formula for finding the vertex: \(\large x = \frac{-(b)}{2a}\) For \(y = x^2 -6x + 4\), you know that the b term is -6 and the a term is 1. Just substitute those terms back to the formula. \(\large x = \frac{-(-6)}{2(1)} = \frac{6}{2} = 3\) Knowing that 3 is the x-coordinate for the vertex of the quadratic, we need to know what the y-coordinate is going to be. So now you need to plug it back into the equation. \(y = (3)^2 - 6(3) + 4\) \(y = 9 - 6(3) + 4\) \(y = 9 - 18 + 4\) \(y = -9 + 4\) \(\color{green}{y = -5}\) The vertex is (3, -5).

OpenStudy (fairegaming):

yeah I know that. I already have that on my paper, but there were a few more steps that I didn't really get to in class

OpenStudy (calculusxy):

Since the a value is positive, we can assume that the parabola will be facing upwards. Therefore, the vertex (3, -5) is the minimum point of the parabola. Now, start to plug some numbers in, but I suggest you to do it in this way: